Circular No. 8121

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2003du IN UGC 9391 Further to IAUC 8120, M. Schwartz and P. R. Holvorcem report the

LOTOSS discovery of an apparent supernova (mag 15.9) on unfiltered CCD images taken on Apr. 22.4 (Tenagra IV 0.36-m telescope) and 23.3 UT (Tenagra II 0.81-m telescope). SN 2003du is located at α = 14^{h}34^{m}35^{s}.80, δ = +59^{o}20′03′′.8 (equinox 2000.0), which is 8′′.8 west and 13′′.5 south of the center of UGC 9391. A Tenagra II image taken on Apr. 7.4 showed nothing at this position (limiting mag ∼ 19.0).

POSSIBLE SUPERNOVA IN UGC 9638 Schwartz and Holvorcem also report the discovery of an object (mag

16.2) on unfiltered CCD images taken on Apr. 22.4 and 23.4 UT. The new star is located at α = 14^{h}58^{m}04^{s}.92, δ = +58^{o}52′49′′.9 (equinox 2000.0), which is 37′′.5 west and 31′′.8 north of the center of UGC 9638. Nothing is visible at this location on Tenagra II images taken on Mar. 14.5 and 17.4 (limiting mag 18.5). The new object is 3′′.0 east and 0′′.7 north of a USNO-A2.0 star of mag 18.2; this star is partially involved with the new object in the Apr. 22.4 and 23.4 images.

SUPERNOVAE 2002kz, 2002la, 2002lb, 2002lc, 2002ld Further to IAUC 8119 and 8120, M. Doi reports five additional appar-

ent supernovae that lack spectroscopic confirmation.

SN  2002 UT  α_{2000}  δ_{2000}  i′  Offset
2002kz Nov. 3.3 2^{h}16^{m}56.^{s}37 − 5^{o}00′57′′.4 24.9 1′′.5 southeast
2002la Nov. 3.3 2 17 45.97 − 4 36 46.2 25.2 0′′.2 west
2002lb Nov. 3.3 2 18 09.01 − 4 54 17.9 25.1 0′′.6 southeast
2002lc Nov. 3.3 2 18 51.59 − 4 47 24.8 25.1 0′′.2 north
2002ld Nov. 3.3 2 20 00.03 − 4 44 20.2 24.3 0′′.5 southeast

Additional magnitudes were obtained on 3–6 nights for each object that yield light curves consistent with those of supernovae; i′-band magnitudes for Dec. 8: SN 2002kz, 25.5; SN 2002la, 25.1; SN 2002lb, 26.2; SN 2002lc, 25.8; SN 2002ld, 27.5.

2003 April 23  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2003 CBAT  Daniel W. E. Green